OpenStudy (anonymous):

Which of the following is the slope between the points (-1, -5) and (7, -3)? 4 1/4 -1/4 -4

OpenStudy (anonymous):

Slope can be found as: \[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

\[y_2 = -3, y_1 = -5, x_2 = 7, x_1 = -1\]

hero (hero):

y = mx + b -5 = -1m + b -3 = 7m + b m - 5 = b -3 - 7m = b b = b m - 5 = -3 - 7m 7m + m = 5 - 3 8m = 2 m = 2/8 m = ?